OpenStudy (anonymous):

A 24 kg gazelle runs away from a cheetah going 18 m/s. Calculate the gazelle's kinetic energy

OpenStudy (anonymous):

the kinetic energy would 5 miles per hour

OpenStudy (anonymous):

How did you get it??

OpenStudy (anonymous):

Kinetic energy is defined as: KE = (1/2) * m * v^2 m is the mass v is the speed

OpenStudy (anonymous):

thank you and how do I get potential energy? @Pompeii00

OpenStudy (anonymous):

Potential energy can be defined a couple ways: (1) If there is constant total energy TE, then potential energy PE is equal to the portion of energy that is not kinetic energy KE. (PE = TE - KE). (2) Potential energy gravitationally is defined as: PE = m * g * h m is the mass g is the gravitational acceleration h is the height If you're asking about the potential energy of the gazelle here, then based on the information given the PE would be 0.

OpenStudy (anonymous):

what is the difference between gravitational energy and potential gravitational energy?

OpenStudy (anonymous):

gravitational energy isn't a type of energy. I was just referring to PE=mgh as specifically being gravitational PE since the height of the object is the reason for the PE. PE can also manifest itself in springs for example. That isn't gravitational PE, just PE (or maybe elastic PE).

OpenStudy (anonymous):

oh okay I have one last question, when your done solving for, for example kinetic energy what would be the units if you put units

OpenStudy (anonymous):

The units of energy is Joules [ J ]. Joules, written in terms of other units, is [N m] or [kg m^2 / s^2] Essentially units of force multiplied by units of distance.

OpenStudy (anonymous):

would the units for potential energy be joules too?

OpenStudy (anonymous):

Yep, all types of energy are measured in Joules.

OpenStudy (anonymous):

thanks this was very helpful!

OpenStudy (anonymous):

No problem.